OpenStudy (wes_is_here):

The coordinates below represent two linear equations. How many solutions does this system of equations have? Line 1 x y –1 6 0 4 Line 2 x y 0 7 1 5 A.0 B.exactly 1 C.exactly 2 D.infinitely many

OpenStudy (zpupster):

infinitely many if the 2 lines were the same line, they would have the same slope exactly 1 if the 2 lines crossed 0 if they were parallel or again, they would have the same slope. so off to our slope formula \[m=\frac{ y2-y1 }{ x2-x1 }\] plugging your first set of points from Line 1 m= 4-6/0-(-1)=? DO THIS FOR LINE 2 then we will compare slopes first.

OpenStudy (wes_is_here):

??

OpenStudy (zpupster):

ok lets take it slower

OpenStudy (wes_is_here):

ok

OpenStudy (zpupster):

the questions is asking about solutions to the lines if we look at the slope of the lines --and graph the lines we can determine the number of solutions it has we do this by using the slope formula: \[m=\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (wes_is_here):

ok

OpenStudy (zpupster):

m= 4-6/0-(-1)=? I did the first one for you. if you look at the table for line 1 y2=4 y1=6 x2=0 x1=-1 i plugged them in the slope formula m= 4-6/0-(-1)=? just preform the math. that will give you the slope for Line 1 the we will do that for Line 2 and compare the slopes.

OpenStudy (wes_is_here):

im still confused

OpenStudy (zpupster):

m= 4-6/0-(-1)=? just preform the math. try this first

OpenStudy (wes_is_here):

i dont know how

OpenStudy (zpupster):

well what is 4-6

OpenStudy (wes_is_here):

2

OpenStudy (zpupster):

no since 6 is bigger than 4 it is??

OpenStudy (wes_is_here):

-2

OpenStudy (zpupster):

yes now do 0-(-1)

OpenStudy (wes_is_here):

-1

OpenStudy (zpupster):

since there are 2 negatives in front of the one it is 0+1=

OpenStudy (wes_is_here):

so its 1

OpenStudy (zpupster):

so our slope is -2/1 0r just -2

OpenStudy (zpupster):

remember Line 1 form your chart? Line 1 y2=4 y1=6 x2=0 x1=-1 Line 2 fill in Line 2 from oyur chart y2= y1= x2= x1=

OpenStudy (wes_is_here):

yeah

OpenStudy (wes_is_here):

?

OpenStudy (wes_is_here):

@zpupster

OpenStudy (zpupster):

sorry had to leave. Line 2 fill in Line 2 from your chart in your question. y2= y1= x2= x1=

OpenStudy (wes_is_here):

??